Grade 11Mechanics

a metre scale weighing 600gm ,is displaced through an angle of 60 degrees in vertical plane as shown.the increase in its potential energy is found to be 0.3 into x,find the value of x

rishika
8 Years agoGrade 11

1 Answer

Rituraj Tiwari
5 Years ago
dear student,
Given,m=500gm=0.6kg
θ=60°
Change in potential energy at the center of the meter stick,
ΔU=Uf​−Ui​
ΔU=(−mghcosθ)−(−mgh)=mgh−mghcos600
ΔU=mgh(1−1/2)=mgh​/2
ΔU=(0.6×10×1/2)*2
ΔU=1.5J
hence the value of x will be 5
thank you